# A new spin on polynomial relations among kappa classes

### v1

@@ 1. Introduction
For any $g\geq 0$, $n\geq 0$, 2g2+n>02g-2+n>0, $d\geq 0$ the Poincaré dual of $(-1)^{g}p_{*}C_{g,n,d}$ is equal to $J_{2g-2+n-d}$ (in $R^{*}(\overline{\mathcal{M}}_{g,n})$).
@@
(5) $\langle\tau_{k_{1}}\dots\tau_{k_{n}}\rangle^{\mathbb{P}^{1},\mathcal{O}(-1),{\rm dec}}_{g}\coloneqq\sum_{d=0}^{\infty}(-1)^{g+d}Q^{d}\int_{C_{g,n,d}}\prod_{i=1}^{n}\psi_{i}^{k_{i}}$
@@
(7) $\omega^{(g)}_{n}=2^{2g-2+n}\sum_{k_{1},\dots,k_{n}}\langle\tau_{k_{1}}\dots\tau_{k_{n}}\rangle^{\mathbb{P}^{1},\mathcal{O}(-1),{\rm dec}}_{g}\prod_{i=1}^{n}d\frac{(2k_{i}-1)!!}{x_{i}^{2k_{i}+1}}+\delta_{g,0}\delta_{n,2}\frac{dx_{1}dx_{2}}{(x_{1}-x_{2})^{2}}.$
@@ 2.1. The formula
We fix integers $n\geq 0$ and $g\geq 0$. We denote by $\mathrm{SRT}_{g,n}(\mathbb{P}^{1},d)$ the set of star rooted trees of genus $g$ with $n$ legs, with edges of total degree $d$, which are the graphs
@@
We have:
(10) $C_{g,n,d}(T)=(-1)^{g+d}\;T^{2g-2+n-d}\\ \times\sum_{\Gamma\in\mathrm{SRT}_{g,n,d}}\frac{(-1)^{|E|}}{|E|!}(B_{\Gamma})_{*}\left(\prod_{i=1}^{|E|}\frac{d_{i}^{d_{i}+1}}{d_{i}!T}\frac{1}{1-d_{i}\frac{\psi_{n+i}}{T}}\otimes\bigotimes_{i=1}^{|E|}\frac{\frac{(-1)^{g_{i}}\lambda_{g_{i}}}{T^{g_{i}}}\Lambda(\frac{1}{T})}{1+d_{i}\frac{\psi_{1}}{T}}\right).$
@@ 2.2. Ancestor correlators: numerical evidence
(12) $C_{g,n}(T,Q)=\sum_{d=1}^{\infty}(-1)^{g+d}Q^{d}p_{*}C_{g,n,d}(T)\in R^{*}(\overline{\mathcal{M}}_{g,n})\otimes_{\mathbb{Q}}\mathbb{Q}[T,Q].$
@@
(13) $\langle\tau_{k_{1}}\dots\tau_{k_{n}}\rangle^{\mathbb{P}^{1},\mathcal{O}(-1),\mathbb{C}^{*},{\rm anc}}_{g}=\int_{\overline{\mathcal{M}}_{g,n}}C_{g,n}(T,Q)\;\psi_{1}^{k_{1}}\dots\psi_{n}^{k_{n}}.$
The general theory implies that this quantity is a polynomial in $T$ and $Q$ of homogeneous degree $d_{\rm tot}=\sum k_{i}-g+1$. In particular, it is identically equal to zero if $d_{\rm tot}<0$. The regular dependence of this expression in $Q$ is seen from the very its form. On the other hand, its regular dependence in $T$ (that is, the fact that the contributions of all summands with $d>d_{\rm tot}$ cancel out), is a manifestation of the localization formula. Below, we compute these correlators in a closed form that allows one to take a limit as $T\to 0$. In the next section, we reproduce the same computations directly for the classes $C_{g,n}(T,Q)$ themselves.
Let us collect these correlators to the potential
(14) $F^{\mathbb{P}^{1},\mathcal{O}(-1),\mathbb{C}^{*},{\rm anc}}(t_{0},t_{1},\dots;\hbar)=\sum_{g,n}\frac{\hbar^{2g-2+n}}{n!}\sum_{k_{1},\dots,k_{n}\geq 0}\langle\tau_{k_{1}}\dots\tau_{k_{n}}\rangle^{\mathbb{P}^{1},\mathcal{O}(-1),\mathbb{C}^{*},{\rm anc}}_{g}t_{k_{1}}\dots t_{k_{n}}.$
@@
with the following reservation: the symbol $\approx$ here and below means that the equality holds up to unstable terms with $2g-2+n\leq 0$.
The exceptional terms not covered by Proposition 2.5 are also known; they are treated separately, see below. Note that the shifts of times $t_{k}$ to finite values in the Kontsevich-Witten potential are well defined for $k\geq 2$ only. In our case the shifts involve $\hbar$, and since the expansions of $\widehat{P}_{0}/\hbar$ and $\widehat{P}_{1}/\hbar$ contain strictly positive powers of $\hbar$ only, the shifts of $t_{0}$ and $t_{1}$ are also well defined if treated in the expansion in $\hbar$. This formula expresses each ancestor correlator as a finite combination of intersection numbers of $\psi$ classes with coefficients in $\mathbb{Q}[T,(T-Q)^{\pm 1}]$. The geometrical origin of the potential implies that the poles at $T=Q$ cancel out and the coefficients of the potential are actually polynomial in $T$ and $Q$. Taking the limit at $T\to 0$ we get
@@
(20) $F^{\mathbb{P}^{1},\mathcal{O}(-1),\mathbb{C}^{*},{\rm anc}}\bigm{|}_{T=0}$ $=F^{\rm KW}\bigl{(}t_{0},t_{1},t_{2}+\tfrac{1!}{\hbar Q^{2}},t_{3}+\tfrac{2!}{\hbar Q^{3}},t_{4}+\tfrac{3!}{\hbar Q^{4}},\dots;-Q\hbar).$
@@
(21) $\langle\tau_{k_{1}}\dots\tau_{k_{n}}\rangle^{\mathbb{P}^{1},\mathcal{O}(-1),\mathbb{C}^{*},{\rm des}}_{g}=\sum_{d=1}^{\infty}(-1)^{g+d}Q^{d}\int_{C_{g,n,d}(T)}\psi_{1}^{k_{1}}\dots\psi_{n}^{k_{n}}$
(the nonequivariant limit $\mathop{\big{\lfloor}_{{T}\to{0}}}$ of this formula is given by (5)).
@@
(22) $(-1)^{g+d}C_{g,n,d}(T)=T^{-d}\sum_{\Gamma\in\mathrm{SRT}_{g,n,d}}\frac{T^{2g_{0}-2+n+|E|}}{|E|!}(-1)^{|E|}\times\\ (B_{\Gamma})_{*}\left(\prod_{i=1}^{|E|}\frac{\frac{d_{i}^{d_{i}-1}}{d_{i}!}}{1-\frac{d_{i}}{T}\psi_{n+i}}\otimes\bigotimes_{i=1}^{|E|}\frac{\bigl{(}\frac{d_{i}}{T}\bigr{)}^{2}(-T)^{g_{i}}\lambda_{g_{i}}\Lambda(\frac{1}{T})}{1+\frac{d_{i}}{T}\psi_{1}}\right).$
@@
(23) $\langle\tau_{k_{1}}\dots\tau_{k_{n}}\rangle^{\mathbb{P}^{1},\mathcal{O}(-1),\mathbb{C}^{*},{\rm dec}}_{g}\approx\sum_{g_{0}=0}^{g}T^{2g_{0}-2+n}\langle\tau_{k_{1}}\dots\tau_{k_{n}}[\hbar^{2(g-g_{0})}]e^{-T\sum_{k=0}^{\infty}P_{k}\tau_{k}}\rangle^{\rm pt}_{g_{0}},$
@@
Here by $\langle\cdot\rangle^{\rm pt}$ we denote the Kontsevich-Witten correlators, that is, those enumerating intersection numbers of just $\psi$ classes over moduli spaces of curves. Again, the symbol $\approx$ means that the equality is not applicable in the cases $2g-2+n\leq 0$ and $n\leq 1$; these cases should be treated separately. Then, the obtained expression for the correlators leads to the corresponding formula for the potential:
@@
$\approx F^{\rm KW}\bigl{(}\sum_{m=0}^{\infty}\tfrac{Q^{m}}{m!}t_{m}-\tfrac{P_{0}}{\hbar},\sum_{m=0}^{\infty}\tfrac{Q^{m}}{m!}t_{1+m}-\tfrac{P_{1}}{\hbar},\dots;T\hbar).$
$=F^{\rm KW}\bigl{(}\sum_{m=0}^{\infty}\tfrac{Q^{m}}{m!}\bigl{(}t_{m}-\tfrac{\widetilde{P}_{m}}{\hbar}\bigr{)},\sum_{m=0}^{\infty}\tfrac{Q^{m}}{m!}\bigl{(}t_{1+m}-\tfrac{\widetilde{P}_{1+m}}{\hbar}\bigr{)},\dots;T\hbar),$
@@
$=\sum_{m=0}^{\infty}\sum_{d=1}^{\infty}\Bigl{(}\frac{Q}{T}\Bigr{)}^{d}\;\frac{d^{d-1}}{d!}\frac{(-Q)^{m}\bigl{(}\frac{d}{T}\bigr{)}^{m}}{m!}\Bigl{(}\frac{d}{T}\Bigr{)}^{k}\biggl{(}1+\sum_{g_{1}=1}^{\infty}\hbar^{2g_{1}}\!\!\!\int\limits_{\overline{\mathcal{M}}_{g_{1},1}}\!\!\!\frac{\bigl{(}\frac{d}{T}\bigr{)}^{2}(-T)^{g_{i}}\lambda_{g_{i}}\Lambda(\frac{1}{T})}{1+\frac{d}{T}\psi_{1}}\biggr{)}$
$=\sum_{d=1}^{\infty}X^{d}\;\frac{d^{d-1}}{d!}\Bigl{(}\frac{d}{T}\Bigr{)}^{k}\biggl{(}1+\sum_{g_{1}=1}^{\infty}\hbar^{2g_{1}}\!\!\!\int\limits_{\overline{\mathcal{M}}_{g_{1},1}}\!\!\!\frac{\bigl{(}\frac{d}{T}\bigr{)}^{2}(-T)^{g_{i}}\lambda_{g_{i}}\Lambda(\frac{1}{T})}{1+\frac{d}{T}\psi_{1}}\biggr{)},$
@@
Using known properties of the Lambert curve $X=\tfrac{Q}{T}e^{-\frac{Q}{T}}$ including the identity $T^{-1}X\frac{d}{dX}=\frac{Q}{T-Q}\frac{d}{dQ}$ we compute
@@
Note that both $F^{\mathbb{P}^{1},\mathcal{O}(-1),\mathbb{C}^{*},{\rm des}}$ and $F^{\mathbb{P}^{1},\mathcal{O}(-1),\mathbb{C}^{*},{\rm anc}}$ are solutions of the KdV integrable hierarchy, therefore can be considered in the framework of reductions of both KP and BKP hierarchies.
@@ 2.3. Materialization
(38) $\qquad+\delta_{n,0}\sum_{a=0}^{g-1}(-1)^{g}T^{g-a}{\lambda_{g}\lambda_{a}}\times$
$\qquad\qquad\biggl{(}\sum_{l=0}^{\infty}(-1)^{l}\bigl{(}\tfrac{V_{l+1}}{T-Q}-V_{1,l}\bigr{)}\,\kappa_{l}+\frac{1}{2}\sum_{k,l=0}^{\infty}(-1)^{k+l}V_{k+1,l+1}b_{*}\bigl{(}(\psi^{\prime})^{k}\otimes(\psi^{\prime\prime})^{l}\bigr{)}\biggr{)}.$
@@
It is convenient to abuse the notation and to think of $p_{*}\sum_{d=1}^{\infty}(-1)^{g+d}Q^{d}C_{g,n,d}(T)$ rather than of $\sum_{d=1}^{\infty}(-1)^{g+d}Q^{d}p_{*}C_{g,n,d}(T)$, though it is not a single push-forward but rather a system of push-forwards that depends on $d$. This allows us to take the sum over all labels $d\colon E\to\mathbb{Z}_{\geq 1}$ for each given pre-stable star rooted tree, with no condition on the sum of $d(e)$ over $e\in E$. For the class (10) or (22) rewritten in this way we assemble the dependence on $Q$ for each edge $e$ of genus $g(e)=h$ into the series

### v2

@@ 1. Introduction
For any $g\geq 0$, $n\geq 0$, 2g2+n>02g-2+n>0, $d\geq 0$ the Poincaré dual of $(-1)^{g+d+1}p_{*}C_{g,n,d}$ is equal to $J_{2g-2+n-d}$ (in $R^{*}(\overline{\mathcal{M}}_{g,n})$).
@@
(5) $\langle\tau_{k_{1}}\dots\tau_{k_{n}}\rangle^{\mathbb{P}^{1},\mathcal{O}(-1),{\rm des}}_{g}\coloneqq\sum_{d=0}^{\infty}(-1)^{g+1}Q^{d}\int_{C_{g,n,d}}\prod_{i=1}^{n}\psi_{i}^{k_{i}}$
@@
(7) $\omega^{(g)}_{n}=2^{2g-2+n}\sum_{k_{1},\dots,k_{n}}\langle\tau_{k_{1}}\dots\tau_{k_{n}}\rangle^{\mathbb{P}^{1},\mathcal{O}(-1),{\rm des}}_{g}\prod_{i=1}^{n}d\frac{(2k_{i}-1)!!}{x_{i}^{2k_{i}+1}}+\delta_{g,0}\delta_{n,2}\frac{dx_{1}dx_{2}}{(x_{1}-x_{2})^{2}}.$
@@ 1.3. Acknowledgments
The authors thank Xavier Blot and Reinier Kramer for very useful discussions.
@@ 2.1. The formula
We fix integers $g,n\geq 0$ and $d\geq 1$. We denote by $\mathrm{SRT}_{g,n}(\mathbb{P}^{1},d)$ the set of star rooted trees of genus $g$ with $n$ legs, with edges of total degree $d$, which are the graphs
@@
For $d\geq 1$ we have:
(10) $C_{g,n,d}(T)=(-1)^{g+1}\;T^{2g-2+n-d}\\ \times\sum_{\Gamma\in\mathrm{SRT}_{g,n,d}}\frac{(-1)^{|E|}}{|E|!}(B_{\Gamma})_{*}\left(\prod_{i=1}^{|E|}\frac{d_{i}^{d_{i}+1}}{d_{i}!T}\frac{1}{1-d_{i}\frac{\psi_{n+i}}{T}}\otimes\bigotimes_{i=1}^{|E|}\frac{\frac{(-1)^{g_{i}}\lambda_{g_{i}}}{T^{g_{i}}}\Lambda(\frac{1}{T})}{1+d_{i}\frac{\psi_{1}}{T}}\right).$
@@ 2.2. Ancestor correlators: numerical evidence
(12) $C_{g,n}(T,Q)\coloneqq\sum_{d=1}^{\infty}(-1)^{g+1}Q^{d}p_{*}C_{g,n,d}(T)\in R^{*}(\overline{\mathcal{M}}_{g,n})\otimes_{\mathbb{Q}}\mathbb{Q}[T,Q].$
@@
(13) $\langle\tau_{k_{1}}\dots\tau_{k_{n}}\rangle^{\mathbb{P}^{1},\mathcal{O}(-1),\mathbb{C}^{*},{\rm anc}}_{g}\coloneqq\int_{\overline{\mathcal{M}}_{g,n}}C_{g,n}(T,Q)\;\psi_{1}^{k_{1}}\dots\psi_{n}^{k_{n}}.$
The general theory implies that this quantity is a polynomial in $T$ and $Q$ of homogeneous degree $d_{\rm tot}=\sum k_{i}-g+1$. In particular, it is identically equal to zero if $d_{\rm tot}<0$. The regular dependence of this expression in $Q$ is seen from the very its form. On the other hand, its regular dependence in $T$ (that is, the fact that the contributions of all summands with $d>d_{\rm tot}$ cancel out), is a manifestation of the localization formula. Below, we compute these correlators in a closed form that allows one to take a limit as $T\to 0$. In the next section, we reproduce the same computations directly for the classes $C_{g,n}(T,Q)$ themselves. When we pass to $T\to 0$ limit, we routinely add the $d=0$ terms to the definition of the ancestor correlators, that is, we define
(14) $\langle\tau_{k_{1}}\dots\tau_{k_{n}}\rangle^{\mathbb{P}^{1},\mathcal{O}(-1),{\rm anc}}_{g}\coloneqq(-1)^{g+1}\int_{\overline{\mathcal{M}}_{g,n}}\sum_{d=0}^{\infty}Q^{d}p_{*}C_{g,n,d}\;\psi_{1}^{k_{1}}\dots\psi_{n}^{k_{n}}.$
Let us collect these correlators to the potentials
(15) $F^{\mathbb{P}^{1},\mathcal{O}(-1),\mathbb{C}^{*},{\rm anc}}(t_{0},t_{1},\dots;\hbar)$ $\coloneqq\sum_{g,n}\frac{\hbar^{2g-2+n}}{n!}\sum_{k_{1},\dots,k_{n}\geq 0}\langle\tau_{k_{1}}\dots\tau_{k_{n}}\rangle^{\mathbb{P}^{1},\mathcal{O}(-1),\mathbb{C}^{*},{\rm anc}}_{g}t_{k_{1}}\dots t_{k_{n}};$
$F^{\mathbb{P}^{1},\mathcal{O}(-1),{\rm anc}}(t_{0},t_{1},\dots;\hbar)$ $\coloneqq\sum_{g,n}\frac{\hbar^{2g-2+n}}{n!}\sum_{k_{1},\dots,k_{n}\geq 0}\langle\tau_{k_{1}}\dots\tau_{k_{n}}\rangle^{\mathbb{P}^{1},\mathcal{O}(-1),{\rm anc}}_{g}t_{k_{1}}\dots t_{k_{n}}.$
@@
with the following reservation: the symbol $\approx$ here and below means that the equality holds up to unstable terms with $2g-2+n\leq 0$, and also the terms with $d=0$ or $n\leq 1$ and arbitrary $g$.
The exceptional terms not covered by Proposition 2.5 are also known; they are treated separately, see below. Note that the shifts of times $t_{k}$ to finite values in the Kontsevich-Witten potential are well defined for $k\geq 2$ only. In our case the shifts involve $\hbar$, and since the expansions of $\widehat{P}_{0}/\hbar$ and $\widehat{P}_{1}/\hbar$ contain strictly positive powers of $\hbar$ only, the shifts of $t_{0}$ and $t_{1}$ are also well defined if treated in the expansion in $\hbar$. This formula expresses each ancestor correlator as a finite combination of intersection numbers of $\psi$ classes with coefficients in $\mathbb{Q}[T,(T-Q)^{\pm 1}]$. The geometrical origin of the potential implies that the poles at $T=Q$ cancel out and the coefficients of the potential are actually polynomial in $T$ and $Q$. Taking the limit at $T\to 0$ we get (after including the $d=0$ terms):
@@
(21) $F^{\mathbb{P}^{1},\mathcal{O}(-1),{\rm anc}}$ $=F^{\rm KW}\bigl{(}t_{0},t_{1},t_{2}+\tfrac{1!}{\hbar Q^{2}},t_{3}+\tfrac{2!}{\hbar Q^{3}},t_{4}+\tfrac{3!}{\hbar Q^{4}},\dots;-Q\hbar).$
@@
Remark 2.6.
Note that formally Equation (21) holds exactly as it is presented, unlike the statement of Proposition 2.5, where we use the symbol $\approx$. So it is formally not a direct corollary and some additional analysis of the $d=0$ and/or $n\leq 1$, and unstable terms is required. We refer for that to Section 2.3 and 2.4.
@@
(22) $\langle\tau_{k_{1}}\dots\tau_{k_{n}}\rangle^{\mathbb{P}^{1},\mathcal{O}(-1),\mathbb{C}^{*},{\rm des}}_{g}\coloneqq\sum_{d=1}^{\infty}(-1)^{g+1}Q^{d}\int_{C_{g,n,d}(T)}\psi_{1}^{k_{1}}\dots\psi_{n}^{k_{n}}$
(the nonequivariant limit $\mathop{\big{\lfloor}_{{T}\to{0}}}$ of this formula is given by (5) up to $d=0$ terms).
@@
(23) $(-1)^{g+1}C_{g,n,d}(T)=T^{-d}\sum_{\Gamma\in\mathrm{SRT}_{g,n,d}}\frac{T^{2g_{0}-2+n+|E|}}{|E|!}(-1)^{|E|}\times\\ (B_{\Gamma})_{*}\left(\prod_{i=1}^{|E|}\frac{\frac{d_{i}^{d_{i}-1}}{d_{i}!}}{1-\frac{d_{i}}{T}\psi_{n+i}}\otimes\bigotimes_{i=1}^{|E|}\frac{\bigl{(}\frac{d_{i}}{T}\bigr{)}^{2}(-T)^{g_{i}}\lambda_{g_{i}}\Lambda(\frac{1}{T})}{1+\frac{d_{i}}{T}\psi_{1}}\right).$
@@
(24) $\langle\tau_{k_{1}}\dots\tau_{k_{n}}\rangle^{\mathbb{P}^{1},\mathcal{O}(-1),\mathbb{C}^{*},{\rm des}}_{g}\\ \approx\sum_{g_{0}=0}^{g}T^{2g_{0}-2+n}\langle\tau_{k_{1}}\dots\tau_{k_{n}}[\hbar^{2(g-g_{0})}]e^{-T\sum_{k=0}^{\infty}P_{k}\tau_{k}}\rangle^{\rm pt}_{g_{0}}-T^{2g-2+n}\langle\tau_{k_{1}}\dots\tau_{k_{n}}\rangle^{\rm pt}_{g},$
@@
Here by $\langle\cdot\rangle^{\rm pt}$ we denote the Kontsevich-Witten correlators, that is, those enumerating intersection numbers of just $\psi$ classes over moduli spaces of curves. Again, the symbol $\approx$ means that the equality is not applicable in the cases $n\leq 1$; these cases should be treated separately. Then, the obtained expression for the correlators leads to the corresponding formula for the potential:
@@
$\approx F^{\rm KW}\left(\sum_{m=0}^{\infty}\tfrac{Q^{m}}{m!}t_{m}-\tfrac{P_{0}}{\hbar},\sum_{m=0}^{\infty}\tfrac{Q^{m}}{m!}t_{1+m}-\tfrac{P_{1}}{\hbar},\dots;T\hbar\right)$
$=F^{\rm KW}\left(\sum_{m=0}^{\infty}\tfrac{Q^{m}}{m!}\bigl{(}t_{m}-\tfrac{\widetilde{P}_{m}}{\hbar}\bigr{)},\sum_{m=0}^{\infty}\tfrac{Q^{m}}{m!}\bigl{(}t_{1+m}-\tfrac{\widetilde{P}_{1+m}}{\hbar}\bigr{)},\dots;T\hbar\right),$
@@
$=\sum_{m=0}^{\infty}\sum_{d=1}^{\infty}\Bigl{(}\frac{Q}{T}\Bigr{)}^{d}\;\frac{d^{d-1}}{d!}\frac{(-Q)^{m}\bigl{(}\frac{d}{T}\bigr{)}^{m}}{m!}\Bigl{(}\frac{d}{T}\Bigr{)}^{k}\biggl{(}1+\sum_{g_{1}=1}^{\infty}\hbar^{2g_{1}}\!\!\!\int\limits_{\overline{\mathcal{M}}_{g_{1},1}}\!\!\!\frac{\bigl{(}\frac{d}{T}\bigr{)}^{2}(-T)^{g_{1}}\lambda_{g_{1}}\Lambda(\frac{1}{T})}{1+\frac{d}{T}\psi_{1}}\biggr{)}$
$=\sum_{d=1}^{\infty}X^{d}\;\frac{d^{d-1}}{d!}\Bigl{(}\frac{d}{T}\Bigr{)}^{k}\biggl{(}1+\sum_{g_{1}=1}^{\infty}\hbar^{2g_{1}}\!\!\!\int\limits_{\overline{\mathcal{M}}_{g_{1},1}}\!\!\!\frac{\bigl{(}\frac{d}{T}\bigr{)}^{2}(-T)^{g_{1}}\lambda_{g_{1}}\Lambda(\frac{1}{T})}{1+\frac{d}{T}\psi_{1}}\biggr{)},$
@@
Using known properties of the Lambert function given in our case by the equation $X=\tfrac{Q}{T}e^{-\frac{Q}{T}}$, including the identity $T^{-1}X\frac{d}{dX}=\frac{Q}{T-Q}\frac{d}{dQ}$, we compute
@@
Note that both $F^{\mathbb{P}^{1},\mathcal{O}(-1),\mathbb{C}^{*},{\rm des}}$ and $F^{\mathbb{P}^{1},\mathcal{O}(-1),\mathbb{C}^{*},{\rm anc}}$, up to controllable corrections, are solutions of the KdV integrable hierarchy, therefore can be considered in the framework of reductions of both KP and BKP hierarchies.
@@ 2.3. Materialization
(39) $\qquad+\delta_{n,0}\sum_{a=0}^{g-1}(-1)^{g}T^{g-a}{\lambda_{g}\lambda_{a}}$
$\qquad\qquad\times\biggl{(}\sum_{l=0}^{\infty}(-1)^{l}\bigl{(}\tfrac{V_{l+1}}{T-Q}-V_{1,l}\bigr{)}\,\kappa_{l}+\frac{1}{2}\sum_{k,l=0}^{\infty}(-1)^{k+l}V_{k+1,l+1}b_{*}\bigl{(}(\psi^{\prime})^{k}\otimes(\psi^{\prime\prime})^{l}\bigr{)}\biggr{)}.$
@@
It is convenient to abuse the notation and to think of $p_{*}\sum_{d=1}^{\infty}(-1)^{g+1}Q^{d}C_{g,n,d}(T)$ rather than of $\sum_{d=1}^{\infty}(-1)^{g+1}Q^{d}p_{*}C_{g,n,d}(T)$, though it is not a single push-forward but rather a system of push-forwards that depends on $d$. This allows us to take the sum over all labels $d\colon E\to\mathbb{Z}_{\geq 1}$ for each given pre-stable star rooted tree, with no condition on the sum of $d(e)$ over $e\in E$. For the class (10) or (23) rewritten in this way we assemble the dependence on $Q$ for each edge $e$ of genus $g(e)=h$ into the series